OpenStudy (anonymous):

Assume that y varies inversely as x. If y = -6 when x = -2, find y when x = 5. A. y = 1.7 B. y = 2.4 C. y = 3

OpenStudy (anonymous):

y = k/x is the inverse variation you must firsst find k -6 * - 2= k = 12 y = 12/ 5 = 2.4